OpenStudy (anonymous):

What is the solution set of the following equation? -8x + x + 15 = -7x + 12 Ø {1/7} {all reals

OpenStudy (stamp):

see attachment

OpenStudy (anonymous):

Which of the following represents 3x - 5y + 10 = 0 written in slope-intercept form? y= -x+ 2 y=x- 2 y=x+ 2

OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)? 2x- 3y- 5 = 0 2x+ 3y- 5 = 0 -2x+ 3y- 5 = 0